# Mixed Berndt-Type Integrals and Generalized Barnes Multiple Zeta Functions

### v2

@@ 6 Evaluations of Mixed Berndt-type Integrals
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left[\sinh x-\sin x\right]\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=\left(-4\right)^{m}\pi^{4m-2}\left[\pi\mathrm{DS}_{4m-2,2,1}\left(2\pi\right)-\left(2m-1\right)S_{4m-3,1}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left[\sinh x+\sin x\right]\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=2\left(-4\right)^{m}\pi^{4m}\left[2mS_{4m-1,1}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2,1}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\left(-1\right)^{m}}{2^{2m-2}}\pi^{4m-2}\left[\left(4m-2\right)S_{4m-3,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m-2,2,1}^{\prime}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\left(-1\right)^{m-1}}{2^{2m-1}}\pi^{4m}\left[4mS_{4m-1,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2,1}^{\prime}\left(2\pi\right)\right].$ (6.142)
@@
Finally, we conclude this section by the following conjecture which is supported by comparing the coefficients in Examples 6.2 and 6.5,
@@
For $m\geq 1$, let $c_{2,m},e_{2,m},d_{1,m},f_{1,m}$ be the numbers appearing in Theorems 6.1 and 6.3. Then the following relations hold
@@
7 Berndt-type Integrals via Barnes Multiple Zeta Functions
In this section, we evaluate mixed Berndt-type integrals using the generalized Barnes multiple zeta function. Bradshaw and Vignat [7] evaluated generalized Berndt-type integrals (1.1) using the (alternating) Barnes’ multiple zeta function. For example, they found that ([7, Prop. 2])
@@ 7 Berndt-type Integrals via Barnes Multiple Zeta Functions
[…]
@@
In particular, Z.P. Bradshaw and C. Vignat [7] gave the integral representation of Barnes multiple zeta function, which appears as [15, Eq. (3.2)].
Proposition 7.1.
Let $\Re(s)>N,\Re(w)>0$, and $\Re(a_{j})>0$ for $j=1,\ldots,N$. Then
$\zeta_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1-e^{-a_{j}u})^{-1}du.$
The alternating version is
$\widetilde{\zeta}_{N}(s,w|a_{1},\ldots,a_{N})=\displaystyle\frac{1}{\Gamma(s)}\displaystyle\int_{0}^{\infty}u^{s-1}e^{-wu}\prod_{j=1}^{N}(1+e^{-a_{j}u})^{-1}du.$
Here we define the generalized Barnes multiple zeta function by
@@
Since $\Re(s)>N$, we can write the gamma function in its integral representation, and the series representation of $\zeta_{N}$ converges absolutely for $\Re(w)>0$ and $\Re(a_{j})>0$. We have
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})\Gamma\left(s\right)$ $=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\displaystyle\frac{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}\displaystyle\int_{0}^{\infty}{x^{s-1}e^{-x}\mathrm{d}x}}\quad$
$=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\displaystyle\int_{0}^{\infty}{\displaystyle\frac{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}x^{s-1}e^{-x}}{(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})^{s}}}\mathrm{d}x}$
$=\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\left(\sigma_{1}\right)^{n_{1}}\cdots\left(\sigma_{N}\right)^{n_{N}}\displaystyle\int_{0}^{\infty}{u^{s-1}e^{-(\omega+n_{1}a_{1}+\cdots+n_{N}a_{N})u}}\mathrm{d}u}.$
The sum can be moved inside by the dominated convergence theorem, namely
$\zeta_{N}(s,\omega|\mathbf{a}_{N};\boldsymbol{\sigma}_{N})\Gamma\left(s\right)$ $=\displaystyle\int_{0}^{\infty}{u^{s-1}}e^{-\omega u}\displaystyle\sum_{n_{1}\geq 0,...,n_{N}\geq 0}{\left(\sigma_{1}e^{-a_{1}u}\right)^{n_{1}}\cdots\left(\sigma_{N}e^{-a_{N}u}\right)^{n_{N}}\mathrm{d}u}$
$=\displaystyle\int_{0}^{\infty}{u^{s-1}}e^{-\omega u}\prod_{j=1}^{N}{\left(1-\sigma_{j}e^{-a_{J}u}\right)^{-1}}\mathrm{d}u.$
Noting that $\Re(a_{j})>0$, the geometric series formula ensures that the last equality holds. ∎
@@
$\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}\zeta_{4}\left(4m+1,3+i^{j}|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$
$\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right)$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}},$
$\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4})$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$
$\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4}\right)}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}}\quad(m\geq 2),$
@@
$\displaystyle\frac{\left(\sinh x\pm\sin x\right)}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}=\displaystyle\frac{e^{x}-e^{-x}\pm e^{ix}\mp e^{-ix}}{4\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)},$
$\displaystyle\frac{\left(\sinh x\pm\sin x\right)}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}=\displaystyle\frac{e^{x}-e^{-x}\pm e^{ix}\mp e^{-ix}}{4\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)},$
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{s-1}e^{-\omega x}}{\prod_{i=1}^{M}{\sinh\left(a_{1}x\right)}\prod_{j=1}^{N}{\cosh\left(b_{j}x\right)}}}\mathrm{d}x$ $=\displaystyle\sum_{\tiny\begin{array}[]{c}n_{1},\cdots,n_{M}\geq 0\\ k_{1},\cdots,k_{N}\geq 0\\ \end{array}}{\displaystyle\frac{2^{M+N}\Gamma\left(s\right)\left(-1\right)^{k_{1}+\cdots+k_{N}}}{\left(w+\displaystyle\sum_{i=1}^{M}{a_{i}+\displaystyle\sum_{i=1}^{N}{b_{i}+}2\left(\displaystyle\sum_{i=1}^{M}{a_{i}n_{i}}+\displaystyle\sum_{i=1}^{N}{b_{i}k_{i}}\right)}\right)^{s}}}$
$=2^{M+N}\Gamma\left(s\right)\zeta_{M+N}\left(s,w+\displaystyle\sum_{i=1}^{M}{a_{i}+}\displaystyle\sum_{i=1}^{N}{b_{i}}|\boldsymbol{c}_{M+N},\boldsymbol{\sigma}_{M+N}\right),$
where $\boldsymbol{c}_{M+N}=\left(2a_{1},\cdots,2a_{M},2b_{1},\cdots,2b_{N}\right),\boldsymbol{\sigma}_{M+N}=\left(\left\{1\right\}^{M},\left\{-1\right\}^{N}\right).$ Thus, we deduce that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sin x+\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(e^{x}-e^{-x}+e^{ix}-e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=\displaystyle\frac{1}{4}\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}e^{-i^{j}x}\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right),$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(e^{x}-e^{-x}-e^{ix}+e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=\displaystyle\frac{1}{4}\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}e^{-i^{j}x}\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}}\zeta_{4}\left(4m+1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right),$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(e^{x}-e^{-x}+e^{ix}-e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\tfrac{j^{2}-5j+2}{2}}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4}),$
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$
$=\displaystyle\frac{1}{4}\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(e^{x}-e^{-x}-e^{ix}+e^{-ix}\right)\mathrm{d}x}{\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)}}$
$=4\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-1\right)^{\lfloor\tfrac{j-1}{2}\rfloor}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4}),$
Finally, using Theorems 6.1 and 6.3 yield the desired evaluations. ∎
Letting $m=1,2$, by Mathematica we can compute the following easily using the formulas in Theorem 7.3:
@@
$\zeta_{4}\left(5,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(5,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(5,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(5,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$\zeta_{4}\left(9,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(9,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(9,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(9,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$\zeta_{4}\left(3,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(3,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(3,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(3,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$\zeta_{4}\left(7,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(7,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(7,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(7,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$\zeta_{4}(5,2|\mathbf{a}_{4})+\zeta_{4}(5,3-i|\mathbf{a}_{4})-\zeta_{4}(5,3+i|\mathbf{a}_{4})-\zeta_{4}(5,4|\mathbf{a}_{4})=\displaystyle\frac{\Gamma^{12}}{786432\pi^{4}}-\displaystyle\frac{\Gamma^{8}}{98304\pi^{2}},$
$\zeta_{4}(9,2|\mathbf{a}_{4})+\zeta_{4}(9,3-i|\mathbf{a}_{4})-\zeta_{4}(9,3+i|\mathbf{a}_{4})-\zeta_{4}(9,4|\mathbf{a}_{4})=\displaystyle\frac{9\Gamma^{16}}{4697620480\pi^{4}}-\displaystyle\frac{11\Gamma^{20}}{112742891520\pi^{6}},$
$\zeta_{4}(7,3+i|\mathbf{a}_{4})+\zeta_{4}(7,2|\mathbf{a}_{4})-\zeta_{4}(7,3-i|\mathbf{a}_{4})-\zeta_{4}(7,4|\mathbf{a}_{4})=\displaystyle\frac{\Gamma^{16}}{75497472\pi^{5}}-\displaystyle\frac{\Gamma^{12}}{5242880\pi^{3}},$

### v4

@@ 5 Evaluations of Mixed Berndt-type Integrals
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left[\sinh x-\sin x\right]\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=\left(-4\right)^{m}\pi^{4m-2}\left[\pi\mathrm{DS}_{4m-2,2}\left(2\pi\right)-\left(2m-1\right)S_{4m-3,1}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left[\sinh x+\sin x\right]\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=2\left(-4\right)^{m}\pi^{4m}\left[2mS_{4m-1,1}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\left(-1\right)^{m}}{2^{2m-2}}\pi^{4m-2}\left[\left(4m-2\right)S_{4m-3,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m-2,2}^{\prime}\left(2\pi\right)\right],$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}=\displaystyle\frac{\left(-1\right)^{m-1}\pi^{4m}}{2^{2m-1}}\left[4mS_{4m-1,1}^{\prime}\left(2\pi\right)-\pi\mathrm{DS}_{4m,2}^{\prime}\left(2\pi\right)\right].$ (5.124)
@@
Finally, we conclude this section by the following conjecture which is supported by comparing the coefficients in Examples 5.3 and 5.6:
@@
For $m\geq 1$, let $c_{2,m},e_{2,m},d_{1,m},f_{1,m}$ be the numbers appearing in Theorems 5.1 and 5.4. Then the following relations hold
@@
Next, we evaluate mixed Berndt-type integrals using the generalized Barnes multiple zeta function. Bradshaw and Vignat [7] evaluated generalized Berndt-type integrals (1.1) using the (alternating) Barnes’ multiple zeta function. For example, they found that ([7, Prop. 2])
@@
where ${\boldsymbol{\sl{s}}}^{n}$ means the string $\boldsymbol{\sl{s}}$} is repeated $n$ times. Obviously, combining this with the results of Xu and Zhao [24] yields closed-form evaluations of the Barnes multiple zeta function, see [7, Cor. 2]. Here for positive real numbers $a_{1},\ldots,a_{N}$, the Barnes multiple zeta function and alternating Barnes multiple zeta function are defined by [3, 4, 7, 11] (${\bf a}_{N}:=\{a_{1},\ldots,a_{N}\}$)
@@
In particular, Z.P. Bradshaw and C. Vignat [7] gave the integral representation of Barnes multiple zeta function, which appears as [17, Eq. (3.2)]. Here we define the generalized Barnes multiple zeta function by
@@
This proof is analogous to that in [7, Prop. 2]. ∎
@@
$\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{i^{j-2}\zeta_{4}\left(4m+1,3+i^{j}|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$
$\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-i\right)^{j-2}}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right)$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m-1}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}},$
$\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-i\right)^{j-2}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4})$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m+4}}{\pi^{2m+2}},$
$\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{i^{j-2}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4}\right)}$ $\in\mathbb{Q}\displaystyle\frac{\Gamma^{8m-4}}{\pi^{2m}}+\mathbb{Q}\displaystyle\frac{\Gamma^{8m}}{\pi^{2m+1}}\quad(m\geq 2),$
@@
$\displaystyle\frac{\left(\sinh x\pm\sin x\right)}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}=\displaystyle\frac{e^{x}-e^{-x}\mp ie^{ix}\pm ie^{-ix}}{4\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\sinh\left(\displaystyle\frac{1-i}{2}x\right)\sinh\left(\displaystyle\frac{1+i}{2}x\right)},$
$\displaystyle\frac{\left(\sinh x\pm\sin x\right)}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}=\displaystyle\frac{e^{x}-e^{-x}\mp ie^{ix}\pm ie^{-ix}}{4\sinh\left(\left(1-i\right)x\right)\sinh\left(\left(1+i\right)x\right)\cosh\left(\displaystyle\frac{1-i}{2}x\right)\cosh\left(\displaystyle\frac{1+i}{2}x\right)}.$
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{s-1}e^{-\omega x}}{\prod_{i=1}^{M}{\sinh\left(a_{1}x\right)}\prod_{j=1}^{N}{\cosh\left(b_{j}x\right)}}}\mathrm{d}x=2^{M+N}\Gamma\left(s\right)\zeta_{M+N}\left(s,w+\displaystyle\sum_{i=1}^{M}{a_{i}+}\displaystyle\sum_{i=1}^{N}{b_{i}}|\boldsymbol{c}_{M+N},\boldsymbol{\sigma}_{M+N}\right),$
where $\boldsymbol{c}_{M+N}=\left(2a_{1},\cdots,2a_{M},2b_{1},\cdots,2b_{N}\right),\boldsymbol{\sigma}_{M+N}=\left(\left\{1\right\}^{M},\left\{-1\right\}^{N}\right).$ Thus, we can easily deduce that
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sin x+\sinh x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=4\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{\left(-i\right)^{j-2}}\zeta_{4}\left(4m-1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right),$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x+\cos x\right]}}$ $=4\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{i^{j-2}}\zeta_{4}\left(4m+1,3+i^{j}|\mathbf{a}_{4},\boldsymbol{\sigma}_{4}\right),$
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m}\left(\sinh x+\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}$ $=4\Gamma\left(4m+1\right)\displaystyle\sum_{j=1}^{4}{\left(-i\right)^{j-2}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4}),$
@@
$\displaystyle\int_{0}^{\infty}{\displaystyle\frac{x^{4m-2}\left(\sinh x-\sin x\right)\mathrm{d}x}{\left[\sinh^{2}x+\sin^{2}x\right]\left[\cosh x-\cos x\right]}}=4\Gamma\left(4m-1\right)\displaystyle\sum_{j=1}^{4}{i^{j-2}}\zeta_{4}(4m+1,3+i^{j}|\mathbf{a}_{4}),$
where $\mathbf{a}_{4}=\left(2-2i,2+2i,1-i,1+i\right)$ and $\boldsymbol{\sigma}_{4}=\left(\left\{1\right\}^{2},\left\{-1\right\}^{2}\right)$. Finally, using Theorems 5.1 and 5.4 yields the desired evaluations. ∎
Using the formulas in Theorem 5.10, we can easily compute the following for $m=1,2$ using Mathematica:
@@
$i\zeta_{4}\left(5,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-i\zeta_{4}\left(5,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(5,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(5,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$i\zeta_{4}\left(9,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-i\zeta_{4}\left(9,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(9,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(9,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$i\zeta_{4}\left(3,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-i\zeta_{4}\left(3,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(3,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(3,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$i\zeta_{4}\left(7,3+i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-i\zeta_{4}\left(7,3-i|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)+\zeta_{4}\left(7,2|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)-\zeta_{4}\left(7,4|\mathbf{a}_{4};\boldsymbol{\sigma}_{4}\right)$
@@
$\zeta_{4}(5,2|\mathbf{a}_{4})-i\zeta_{4}(5,3-i|\mathbf{a}_{4})+i\zeta_{4}(5,3+i|\mathbf{a}_{4})-\zeta_{4}(5,4|\mathbf{a}_{4})=\displaystyle\frac{\Gamma^{12}}{786432\pi^{4}}-\displaystyle\frac{\Gamma^{8}}{98304\pi^{2}},$
$\zeta_{4}(9,2|\mathbf{a}_{4})-i\zeta_{4}(9,3-i|\mathbf{a}_{4})+i\zeta_{4}(9,3+i|\mathbf{a}_{4})-\zeta_{4}(9,4|\mathbf{a}_{4})=\displaystyle\frac{9\Gamma^{16}}{4697620480\pi^{4}}-\displaystyle\frac{11\Gamma^{20}}{112742891520\pi^{6}},$
$i\zeta_{4}(7,3-i|\mathbf{a}_{4})-i\zeta_{4}(7,3+i|\mathbf{a}_{4})+\zeta_{4}(7,2|\mathbf{a}_{4})-\zeta_{4}(7,4|\mathbf{a}_{4})=\displaystyle\frac{\Gamma^{16}}{75497472\pi^{5}}-\displaystyle\frac{\Gamma^{12}}{5242880\pi^{3}},$
@@
Acknowledgements: The author would like to thank the referee for his/her careful reading of the paper and many helpful comments.